OpenStudy (anonymous):

How many years will it take for a $400 investment to double in value if interest is compounded annually at 6% ?

OpenStudy (anonymous):

P(1+R/n)^nt 400(1+.06)^t=800 (1.06)^t=2

OpenStudy (anonymous):

I got that

OpenStudy (anonymous):

I did the work, I just want to check my answer

OpenStudy (anonymous):

sure what'd you get?

OpenStudy (anonymous):

198.26

OpenStudy (anonymous):

no, it will take much less than 198.26 years

OpenStudy (anonymous):

12 years

OpenStudy (anonymous):

Gahhhhhh!! I dont know what to do then :(

OpenStudy (anonymous):

(1.06)^t=2 log 1.06^t =log 2 t log 1.06=log 2 t=log 2 / log 1.06

OpenStudy (anonymous):

use the last equation imran gve u in his first post then take natural log of both sides and solve for t

OpenStudy (anonymous):

Isnt it (1.06)^.06t= 2??

OpenStudy (anonymous):

Show me what you did

OpenStudy (anonymous):

Oh wait! Nevermind!!!!

OpenStudy (anonymous):

Sorry! Haha

OpenStudy (anonymous):

I have another problem.... And i think it's the same answer, can you take a look at it please?

OpenStudy (anonymous):

yes sure,post it

OpenStudy (anonymous):

How many years will it take for a $160 investment to double in value if interest is compounded annually at 6%?

OpenStudy (anonymous):

it would be same answer because final amount is twice as much as initial amount. so the equation will end up being same as above (1.06)^t=2 so same number

OpenStudy (anonymous):

Okay that's what i thought, thank you :))))